OpenStudy (anonymous):

cell phone call costs $0.07 per minute for the first 25 minutes and $0.10 per minute for each additional minute.how much would a 47 minute call cost

OpenStudy (anonymous):

47 minutes is broken down into 25 minutes + (47-25=22)minutes. The first 25 are charged at 7c/min for a total of 7 x 25 = 175c The next 22 minutes are charged at 10c/min for a total of 10 x 22 = 220c The total cost is therefore, (175 + 220)c = 395c which is $3.95

OpenStudy (anonymous):

0.07*25+0.10*(47-25)=1.75+2.2=3.95

OpenStudy (anonymous):

Peter is distributing pamphlets about dog care and samples of dog biscuits.The dog biscuits come in packages of 12 and the pamphlets are in paackages of 20.What is the smallest number of samples and pamphlets he needs to distribute without having any left over? How many packages of dog biscuits and pamphlets will peter need?

OpenStudy (anonymous):

Let the number of dog biscuit packages be D and the number of pamphlet packages, P. The number of dog biscuits distributed will then be 12D (12 biscuits per package times the number of packages) and the number of pamphlets will be 20P by similar reasoning. We want a situation where\[12D=20P\]since that means the number of dog biscuits distributed will equal the number of pamphlets. We can simplify that expression as\[3D=5P\]after dividing by a common factor of 4.

OpenStudy (anonymous):

Well, if D=5 and P=3, the equation will be equal.

OpenStudy (anonymous):

Also, for all integers D less than 5, we cannot find an integer for P that will make the statement true, so D=5 is the minimum we can have, and therefore, P=3 the minimum for pamphlet packages.

OpenStudy (anonymous):

5 dog biscuit packages; 3 pamphlet packages.

OpenStudy (anonymous):

How many packages of d and p will peter need though?

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

2

OpenStudy (anonymous):

No...I wrote it out...5 dog biscuit packages and 3 pamphlet packages.

OpenStudy (anonymous):

so 5 packages of dog biscuits and 3 packages of pamphlet packages?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

what is the smallest number of samples and pamphlets he needs to distribute without having any left over??????????

OpenStudy (anonymous):

He has to use 5 dog biscuit samples (each has 12 biscuits, which will mean he'll get through 60 biscuits) and 3 pamphlet packages (each has 20 pamphlets, which will mean he'll get through 60 pamphlets). 60 biscuits, 60 pamphlets = no waste.

OpenStudy (anonymous):

I dont understand????? Im confusedd

OpenStudy (anonymous):

Peter is distributing pamphlets about dog care and samples of dog biscuits.The dog biscuits come in packages of 12 and the pamphlets are in paackages of 20.What is the smallest number of samples and pamphlets he needs to distribute without having any left over? A:60 How many packages of dog biscuits and pamphlets will peter need? A: D=5, P=3.

OpenStudy (anonymous):

Thers a: And B: the a question is what is the smallest number of samples and pamphlets he needs to distribute without having any left over?????????? And the b problem is How many packages of dog biscuits and pamphlets will peter need?

OpenStudy (anonymous):

So can you write like a:=..... And b = ...........

OpenStudy (anonymous):

a) 60 b) 5 packages of dog biscuits, 3 packages of pamphlets.

OpenStudy (anonymous):

Ok thank you verryyyy much

OpenStudy (anonymous):

Okay, kenzielolo...you should become a fan after that!

OpenStudy (anonymous):

I will Thank you sorry i just did not understand it ? :)

OpenStudy (anonymous):

You don't have to apologize for not understanding something. That's why you came here :)

OpenStudy (anonymous):

yupp!!

OpenStudy (anonymous):

You should read through the explanation and try to learn it :)

OpenStudy (anonymous):

I struggle in math!

OpenStudy (anonymous):

But thank You soo Much!!

OpenStudy (anonymous):

For helping Me!

OpenStudy (anonymous):

It's no problem. You a fan yet..? ;p

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

welcomee

OpenStudy (anonymous):

Don't give up on math, even if it's hard...and come back if you need more help. Bye :)

OpenStudy (anonymous):

bye